OpenStudy (anonymous):

find points of intersection for the curves r=cot(x) and r=2cos(x)

OpenStudy (anonymous):

i got : \[x = \frac{\pi}{6}, \frac{5\pi}{6}\]

OpenStudy (anonymous):

thanks that's what i got also. i guess the answer key is wrong

OpenStudy (anonymous):

bah i forgot the obvious answers <.< the statement is also true when cos(x) = 0. That gives us two more answers: \[x = \frac{\pi}{2}, \frac{3\pi}{2}\]

OpenStudy (anonymous):

I guess i should have solved it like this: \[\frac{\cos(x)}{\sin(x)} = 2\cos(x) \Leftrightarrow \cos(x) = 2\sin(x)\cos(x) \Leftrightarrow 0 = 2\sin(x)\cos(x)-\cos(x)\] \[0 = \cos(x)(2\sin(x)-1)\] So then you can see that either cos(x) = 0, or 2sin(x)-1 = 0, and that gives you all four answers.